OpenStudy (anonymous):

solve by substitution: y=2x + 2 & 4x + y = -4

OpenStudy (johnweldon1993):

so you have y=2x + 2 4x + y = -4 see how that first equation tells you what y equals? so substitute that into the second equation for y 4x + 2x + 2 = -4 now solve for x...

OpenStudy (johnweldon1993):

so 4x + 2x + 2 = -4 6x + 2 = -4 -2 -2 6x = -6 x = -1 plug this into your first equation to solve for y